OpenStudy (twistrose):

What is the product in lowest terms? - 5/12*8/13 -20/39 -10/39 -5/39 3/25

OpenStudy (faiqraees):

multiply the numerator with numerator and denominator with denominator

OpenStudy (twistrose):

i dont knoe how to multiply negatives

OpenStudy (faiqraees):

Okay so just some rules whenever you see a negative sign with a number think of it as a negative 1 is being multiplied with it. Like -5 = -1x5 = -5 In the same way -5 x 2 = -1 x 5 x 2 = -1 x 10 = -10 In the same way -5 x -2 = -1 x 5 x -1 x 2 = -1 x -1 x 5 x 2 = 1 x 5 x 2 = 1 x 10 = 10 In a nutshell, negative multiplied negative will give positive number and negative multiplied by positive will give a negative number

OpenStudy (faiqraees):

Got it?

OpenStudy (twistrose):

yeah i guess

OpenStudy (twistrose):

so -5*8 will be....-40>?

OpenStudy (faiqraees):

perfect

OpenStudy (twistrose):

ok so would the denominator of the first be negative as well right

OpenStudy (faiqraees):

no

OpenStudy (faiqraees):

think of it like - 5/12*8/13 \[\large\rm \frac{-5}{12} * \frac{8}{13} = \frac{-5*8}{12*13} \]

OpenStudy (twistrose):

oh cause it had the negative sign at the fraction bar

OpenStudy (faiqraees):

No the negative sign will be with the whole fraction

OpenStudy (twistrose):

oh ok so its normal multiplication

OpenStudy (faiqraees):

Yes

OpenStudy (twistrose):

so its -40/156

OpenStudy (faiqraees):

Now simplify the fraction

OpenStudy (twistrose):

how do you reduce a negative

OpenStudy (faiqraees):

the same way you reduce a positive but just keep the negative sign there Like -15 / 5 = -3

OpenStudy (twistrose):

ok well how do you reduce this fraction

OpenStudy (faiqraees):

Start by dividing both the numerator and denominator by 2

OpenStudy (twistrose):

ok so -20/78

OpenStudy (faiqraees):

Divide by 2 again

OpenStudy (twistrose):

so -10/39

OpenStudy (twistrose):

thanks